OpenStudy (anonymous):

factor the four-term polynomial by grouping. 9q^2=2pq-9q+2p=

OpenStudy (ajprincess):

9q^2+9q=2pq+2p 9q(q+1)=2p(q+1) 9q(q+1)-2p(q+1)=0 (9q-2p)(q+1)=0